A man walks for some distance at 8 km/h, and for an equal distance at 5 km/h .The total time he takes is 3 hours and 15 minutes. Find the distance he had walked.
Answer by oscargut(2103)

Let d = distance walked (total)
Total time = (d/2)/8+ (d/2)/5 = 3.25
d/16 + d/10 = 3.25
(5d+8d)/80=3.25
13d = 260
d = 20
Answer: 20 km
If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com